OpenStudy (anonymous):

Can someone help me with the concepts of a 3D matrix?

OpenStudy (anonymous):

I'm trying to get it into a 2D... for example say I have a 100 x 200 x 3 3D matrix, where 100x200 is the width and height and 3 is just the three colors it can display. I want to multiply all the contents of the 100x200 by the three colors so that its like

OpenStudy (anonymous):

\[\left[\begin{matrix}a & b &c \\ c & d &e... \\ f & g& h ,,, \\ i & j & k ...\end{matrix}\right]\ \left[\begin{matrix} R \\ G \\ B \end{matrix}\right]\]

OpenStudy (anonymous):

Is that even possible?

OpenStudy (anonymous):

No right because they need to have one common length

OpenStudy (anonymous):

Since the last part of the matrix with the size 3 doesn't match up with the height or width of the matrix how can I do matrix multiplication